Grade 12th passAlgebra

The cordinates of the extremities of one diagonal of a square are (1,1) and (-2,-1).find the co ordinates of its other vertices .
Plz help to solve with best approch

Manish kumar
6 Years agoGrade 12th pass

1 Answer

Aditya Gupta
ApprovedApproved Tutor Answer6 Years ago
first find mid point
O= ( – ½, 0)
slope of diag= 2/3
so slope of other diag= – 3/2
hence eqn of other diag: 2(y – 0)= – 3(x+ ½ )........(1)
now, edges are at 90 deg. so m1*m2= – 1
[(y – 1)/(x – 1)]*[(y+1)/(x+2)]= – 1......(2)
solving (1) and (2),
(x, y)= (1/2, – 3/2) or ( – 3/2, 3/2) ---------- these are the reqd vertices
kindly approve :)